OpenStudy (anonymous):

Graph y= 1/2x^2 - x - 6

OpenStudy (anonymous):

type in your question in wolfarmalpha.com you will get the graph

OpenStudy (anonymous):

if u have a graphing calc u can just hit y= and plug it in then hit graph and it will graph it for u

OpenStudy (anonymous):

y= 1/2(x^2 -2 x - 12)=1/2(x-1)^2 -12 - 1/2 (y+ 25/2)=1/2(x-1)^2

OpenStudy (anonymous):

does it help u?

OpenStudy (anonymous):